Angela needs 20 quarts of 50% antifreeze in her radiator. She plans to obtain this by mixing some pure antifreeze(100%) with a appopriate amount of a 40% antifreeze soluation. how many quarts of each should she use?
every time I do this I come up short when I back check. I would like to Thank you in advance for any help

Answer by scott8148(6628)

let x=(amount of pure antifreeze), so 20-x=(amount of 40% antifreeze)

x(100%)+(20-x)(40%)=20(50%)